OpenStudy (faman39):

Which of the following forms a proportion with 10/14 ? a. 21/15 b. 2/7 c. 25/35 d. 7/5

OpenStudy (anonymous):

b

Directrix (directrix):

c. 25/35 25/35 = 10/14 Cross multiply to get 350 = 350